Please Help me solve this problem.
Maylen is 20 years older than her daughter Hazel. In ten years time she will be twice as old as her daughter. How old are Maylene and Hazel ?
Thanks ~

Answer by mananth(16946)

let hazel be x years
maylene = x+20
...
ten years later their ages will be
hazel = x+10
Maylene = x+20+10=x+30
x+30=2(x+10)
x+30 = 2x+20
2x-x=20-10
x=10 hazels age
Maylene = x+20
Maylene will be 30 years
...
m.ananth@hotmail.ca